# Gas Transfer Between the Inner 3-kpc Disk and the Galactic Central Molecular Zone

### v1

@@ 3 Overall Dynamics of the Gas Flows towards the CMZ
Using new data from the MWISP and supplements from other CO surveys (e.g., Dame et al., 2001; Tokuyama et al., 2019; Eden et al., 2020; Schuller et al., 2021), we construct a sketch to illustrate the overall dynamical nature and details of molecular gas flows towards the CMZ. The upper panel of Figure 3 displays an edge-on view of the spatial distribution and motion characteristics of the overall gas flows from the perspective of the observer, while the lower panel shows gas flows dominated by the Galactic bar’s gravitational potential on a face-on view. For the edge-on view, note that the gas inflows extend to in the positive longitude (i.e., the near dust lane DL1, see Su et al., 2024) and in the negative longitude for the far dust lane by assuming its similar length and inclination angle of (between the bar major axis and the Sun-GC) as DL1. The realistic scenario may be much more complex than what we depict here, but this sketch helps us get a comprehensive picture of the entire gas dynamics process towards the CMZ.
@@
Many of the physical processes in the CMZ, including gas dynamics, have been studied in simulations, in which gas is compressed and shocked as a result of the shift from x1 orbit family to x2 orbit family (e.g., Athanassoula, 1992; Englmaier & Gerhard, 1999; Maciejewski et al., 2002; Regan & Teuben, 2004; Sheth et al., 2005; Rodriguez-Fernandez & Combes, 2008; Combes et al., 2014; Sormani et al., 2015). As seen in Figure 3, our observations align well with the model in which the Galactic bar potential (1) channels a substantial amount of molecular gas from the inner Galactic disk at 3.2–3.4 kpc towards the CMZ (i.e., DL1 and the associated kinematic features of CO gas), (2) generates noncircular motions and large-scale shock structures along the near dust lane (e.g., a series of bow shocks revealed by CO gas; Su et al., 2024), (3) governs and regulates the formation and evolution of the CMZ through the overshoot mechanism (i.e., gas flow collision/accumulation near G3/G5/G7 and multiple inflows DL2/DL3 with different trajectories towards the CMZ in Section 2; also see details in the following discussion), and (4) have important impacts on the gas distribution, properties, and the subsequent star formation within the inner Galaxy (e.g., the 3 kpc arm, the thickness of the gas disk, and the star formation rate therein, etc.). As another good example of gas influenced by the non-axisymmetric bar potential, we will further discuss the tilted bar lanes traced by CO flows in Section 4.
@@
The Galactic bar systematically transports a massive inflow towards the CMZ. However, this is only a necessary but not a sufficient condition for the accumulation of gas in the CMZ. Another key question is how to effectively remove the large amount of angular momentum of the gas located a few kpc away, so that the gas can settle continuously into the region within the central few hundred parsecs of the Milky Way. Next, we will focus on how the overshoot effect plays a critical role in regulating the redistribution of gas and its angular momentum, as robustly evidenced by several key kinematic characteristics and dynamical effects observed in the new CO flows towards the CMZ.
@@
In Section 2, we undoubtedly confirmed that G3 originates in violent collisions between the inflowing gas of DL2/DL3 and the overshoot gas from the far dust lane (the G3 case for x1 orbits intersect with x2 orbits; Figure 2b). In this way, G3 is located near the outskirts of the CMZ (i.e., at 0.5 kpc). Meanwhile, gas flow collisions also occur in regions where the overshoot gas changes its trajectory and passes through the bar channel due to varying torques. When the overshoot gas enters the bar channel, collisions between gas flows are inevitable at the intersections of different eccentric orbits (the G5 and G7 cases for gas flow interactions in bar channel; Figures 2a and 2c). Therefore, collisions between gas flows are very common in extremely crowding orbits, especially in the region where the x1 orbits of gas intersects with the x2 orbits in non-axisymmetric gravitational potential (Figure 3).
@@
The starting points of DL2 and DL3 are located exactly around the regions of G7 and G5, respectively. And the gas mass in the intermediate samples is one order of magnitude higher than that in the overshoot gas, confirming significant accumulation of the overshoot gas within the bar channel at these certain locations (e.g., near the G7, G5, and G3 regions). The consistency may indicate that, due to the dynamics of the rotating bar, a significant accumulation of overshoot gas occurs in some special regions (i.e., near G5 at and G7 at by adopting the bar length of 3.2–3.4 kpc; see Su et al., 2024). In addition, considerable CO emission in DL1 is also enhanced significantly near these regions ( and in Figure 2c; also see Figure 1 in Su et al., 2024). These regions are located near the apocentre of the overshoot flows (Figure 3). The interesting feature needs to be further explored.
@@
Importantly, the measured line-of-sight velocities of the DL2 and DL3 inflows towards the CMZ are 1.5–8 times smaller than those of the DL1 at the similar (see Figure 1; also refer to the geometry proposed in Su et al., 2024). The observations clearly indicate that the tangential velocity of gas rotating clockwise around the GC has significantly decreased, and the orbital angular momentum of overshoot gas flows have been reduced by approximately 0.2–0.7 dex after being captured and re-accelerated by the bar potential. These features are consistent with the fact that there is substantial exchange of mass and angular momentum as overshoot gas changes from approaching flows to receding flows at some locations (e.g., the G7 region at 2.2 kpc, the G5 region at 1.6 kpc, and the G3 region at 0.5 kpc; see Figures 2 and 3). Notably, our observations provide the first clear evidence that the overshoot gas, after being captured by the bar potential during its passage through the bar channel, continues to flow towards the CMZ with significantly reduced angular momentum.
@@
The gas mass in DL1 and that associated with the overshoot effect are and (Table 1), respectively. The result indicates that most of the accretion gas does not enter the CMZ directly, but rather overshoots into the bar channel on the other side. Based on our observations, the current inflow efficiency can be easily obtained to be 1/3 on average, which agrees well with the simulation (e.g., Hatchfield et al., 2021). Considering the importance of the overshoot effect (e.g., Maciejewski et al., 2002; Sormani & Barnes, 2019; Hatchfield et al., 2021), we estimate an average inflow rate as . Here, 1/3 is the observed inflow efficiency, while 2 corresponds to both the near and far bars. And 15 Myr comes from the gas kinematics of the DL2 and DL3 inflows by adopting an average inflow velocity of (see DL2 and DL3 in Figure 1). Note that the instantaneous accretion rate can be several times higher than the average value. For example, it can easily cause the accretion rate to increase by nearly five times as several MCs (e.g., with the mass of and the size of pc) with low angular momentum (and thus higher inflow efficiency than 1/3) settle into the CMZ within 0.1 Myr. The gas accumulated in G3/G5/G7 regions serves as a potential reservoir to regulate the instantaneous gas inflow rate.
@@
Again, not all bar-driven gas flows directly enter the CMZ, instead, they take approximately 3 orbital periods to settle into the CMZ (i.e., the observed small inflow efficiency of 1/3). Through this process, the gas is subjected to intense gravitational torques from the bar, causing it to be stretched, distorted, shredded, torn apart, and fragmented (Figure 2c). Apparently, the overshoot effect greatly lengthens the action time of various dynamical processes before the accretion gas enters the CMZ. Throughout the entire dynamical process, the angular momentum of the gas changes and redistributes due to the combined effects such as torque interactions from the bar, inelastic cloud-cloud collisions, gas viscosity, and various instabilities. The shock dissipation generated during the sufficient dynamics will redistribute angular momentum outward. This enables the distorted gas flows to efficiently transport material from kpc scales to regions within a few hundred parsecs of the GC, while also supporting the scenario of the periodic migration of gas towards the CMZ.
@@
Briefly, the confirmation of the overshoot gas, together with frequent collisions caused by intersecting flows, provides detailed observational evidence to explain the dynamics of gas flows towards the CMZ. The overshoot gas with lower angular momentum gradually converges into the bar channel, then flowing towards the CMZ again (see DL2 and DL3 in Figure 3). Because of the overshoot effect and frequent collisions, the accretion gas is able to efficiently transfer its excess angular momentum outward through various dynamical processes, and then causing the gas to gradually flow towards the CMZ. Here, the gravitational torque of the robust rotating Galactic bar makes a large contribution in the process of redistributing the gas and removing angular momentum from the molecular gas that eventually settles into the CMZ. Due to the interaction of various torques with viscous gas and the frequent collisions between gas flows, at the same time, the strong DISSIPATION process is also essential for the effective outward transfer of gas angular momentum, which in turn remarkably regulates the gas flows dynamics and gas accumulation in the CMZ over the secular evolution of the Galaxy.
@@ 4 Large-scale Tilted Gas Flows and The Twisted Structure in the CMZ
The tilted dust lane has been discussed in many literatures (e.g., Liszt, 2006, 2008; Rodriguez-Fernandez et al., 2006; Marshall et al., 2008; Sormani & Barnes, 2019; Tress et al., 2020; Akhter et al., 2021). We also confirm that the near dust lane DL1 with a length of –3.4 kpc exhibits distorted features during the large-scale gas accretion driven by the Galactic bar. Additionally, DL1 is distinctly located below the Galactic plane, which is exactly the opposite of the gas flow of the far dust lane above the plane (see Figure 3).
@@
Our previous studies showed that the slightly tilted mid-plane to the IAU-defined plane (i.e., , see Figure 7 in Su et al., 2016) is caused by the Sun’s height above the physical mid-plane of the Milky Way (i.e., 15–17 pc, see Su et al., 2016, 2019, 2021). However, the tilted large-scale gas flows towards the CMZ cannot be explained by the reason (e.g., see discussions in Bland-Hawthorn & Gerhard, 2016) because the maximum deviation of the CO gas structure reaches 170 pc (or at ) from the Galactic plane of . The deviation value is about 2 times larger than the thickness of the thin CO disk (see Table 3 in Su et al., 2021). Considering the smaller thickness of the molecular gas disk in the inner 3–4 kpc region, this deviation from the plane is even more severe. We attribute the large-scale gas flows to a true tilted feature relative to the Galactic mid-plane. Based on our CO data and model (Su et al., 2024), the tilted angle of the bar lanes relative to the Galactic mid-plane in is estimated to be , which roughly agrees with the simulations (e.g., in Tress et al., 2020). Indeed, the gas structure revealed by the H i emission exhibits a more extreme tilted characteristic (i.e., or 350 pc for the atomic gas with the most negative velocity; see Burton & Liszt, 1983).
@@
The vertical distribution of gas (e.g., the tilted and warped gas flows) is modulated by the strong gravitational torque generated by the rotating bar and various dissipative forces. In our observation, the bar gravitational can effectively trap the gas flows into the dust lane (Figure 3). The gas on different orbits is perturbed by the non-axisymmetric gravitational potential, leading it to undergo non-circular motion around the GC and oscillations relative to the bar. Due to the varying torque exerted by the bar’s potential on the gas, gas flowing along different trajectories and inclinations will move with varying precession rates until it eventually settles into the bar’s potential, aligns with the bar’s major axis, and then rotates around the GC with the bar. These torques also induce misalignments across different parts of the bar lanes. When gas flows on different orbits are subjected to varying torques, they develop distorted or warped structures as observed in CO data. In the process, the angular momentum of the overshoot flows is also transferred to the bar through the dynamic friction. Alternatively, the distortion of the bar lanes caused by resonant trapping of x1 orbits (e.g., Łokas, 2019) may also influence the distribution and motion of the gas. The ongoing interaction between the gas flows and the bar itself accumulates over time, continuously affecting the bar’s structure and motion as it gains mass and angular momentum by exerting torque on the gas.
@@
The above processes are accompanied by complex dynamical interactions that alter the angular momentum and motion states of the gas flows. We confirmed that collisions between gas flows are frequent during the dynamical processes in G3, G5, and G7 regions. On the one hand, the non-axisymmetric bar decouples the gas inflows from their initial dynamical state, resulting in rapid assimilation of the accretion gas into the bar potential (e.g., DL2 and DL3). On the other hand, the frequent collisions between gas flows (G3 and G5), combined with the tangential shear of the tidal forces exerted by the non-axisymmetric bar, result in significant dissipation in the twisted gas flows. Indeed, the typical timescale of these dissipative processes, which are driven by shocks and strong turbulence, may be comparable to the short dynamical timescales associated with gas flow collisions. During the dynamical process of overshoot gas settling into the Galactic bar potential, the dissipation from collisions between misaligned gas flows, together with gas viscosity and the effects of various torques, serve as an efficient mechanism for the outward transfer of angular momentum of gas, and thus facilitating their gradual migration towards the CMZ.
@@
The scenario is strongly supported by our observations that the overshoot gas is not fully relaxed into the Galactic bar potential. That is, the overshoot gas from the far dust lane exhibits a more tilted structure (i.e., 1.5 times larger; see from to in Figure 2c) than the near dust lane DL1 below the Galactic plane, while a portion of the gas passing through the bar channel even extends above the Galactic plane (i.e., DL2 is above and almost parallel to the Galactic plane). The distorted or warped gas structures in observations may represent the oscillatory characteristics of gas flows perpendicular to the Galactic bar as it moves and relaxes in the bar potential.
@@
Meanwhile, the motion state of the self-consistent bar evolves over time through the exchange of angular momentum between the bar itself and the accreted material. As a result, the bar gains angular momentum from the captured gas and loses it to the halo and the outer stellar disk (e.g., Queiroz et al., 2021; Wylie et al., 2022; Dillamore et al., 2023), leading to a gradual change in the state of the bar (e.g., the morphology, structure, and kinematics; see Chiba et al., 2021; Li et al., 2016; Clarke & Gerhard, 2022; Lucey et al., 2023; Vislosky et al., 2024). We suggest that the non-axisymmetric bar is essential for regulating the redistribution of accreted gas and its angular momentum transfer. This self-regulating process may stably maintain large-scale gas accretion and gas recirculation over a considerable timescales in the presence of the bar (e.g., 4–8 Gyr, see Bovy et al., 2019; Nepal et al., 2024).
@@
The accumulated gas in the CMZ inherits the information of the gas supplied by the large-scale inflows, which includes the fresh gas from the disk at 3–4 kpc (with higher angular momentum but lower accretion rate) and the recycling gas from the overshoot effect (with lower angular momentum but higher accretion rate). The moving direction of the large-scale inflows (DL1) is consistent well with the trend of the structure at the far side of the CMZ (see the red solid arrow in the CMZ of Figure 3). That is, gas along the DL1 moves from the negative latitude at towards the positive latitude at , and then back down towards the negative latitude (e.g., see Molinari et al., 2011; Veena et al., 2024; Walker et al., 2024; Lipman et al., 2024; Karoly et al., 2025).
@@
Simultaneously, the inflowing gas along the far dust lane moves from the positive latitude to the negative latitude at , and then towards the positive latitude at , and finally back down towards the negative latitude again (i.e., the near side of the CMZ, see the blue solid arrow in the CMZ of Figure 3). Similar to the previous discussion, when the gas flows pass near the x2 orbit, not all gas directly enters the CMZ, instead, some gas experiences the overshoot effect because of its higher angular momentum. Our data have clearly demonstrated that the overshoot gas flew past the near side of the CMZ and then collided violently with the inflowing gas from the DL2 and DL3 (i.e., see the G3 region in Figure 2b; also see clouds studied by Busch et al., 2022).
@@
We thus propose that the twisted feature of the CMZ (e.g., Molinari et al., 2011; Kruijssen et al., 2015; Ridley et al., 2017) arises from the combined effects of the large-scale tilted gas flows in the dust lanes and the locally distorted gas structure dominated by the strong gravitational potential within a few hundred parsecs of GC (i.e., the nuclear stellar disc; see Sormani et al., 2022). The orbit periods of the twisted structure of the CMZ are about 2–4 Myr (Molinari et al., 2011; Kruijssen et al., 2015), suggesting a rapid precession of the gas under the influence of external non-axisymmetric gravitational potential and the large-scale tilted gas inflows (e.g., Tress et al., 2020). The timescale of several million years mentioned above is consistent with the dynamical time (2–3 Myr) of gas inflow from the G3 region into the CMZ, indicating that external gas input significantly shapes the geometric configuration and dynamic properties of the CMZ (e.g., local gas accumulation, twisted structure, and rapid precession). Recent observations have revealed that the molecular circum-nuclear disks with different kinematics are misaligned with the large-scale gas structure (Combes et al., 2019; Ruffa et al., 2019, 2020; Papachristou et al., 2023), indicating that highly warped gas structures are likely prevalent in the central region of galaxies.

### v2

@@ 3 Overall Dynamics of the Gas Flows toward the CMZ
Using new data from the MWISP and supplements from other CO surveys (e.g., Dame et al., 2001; Tokuyama et al., 2019; Eden et al., 2020; Schuller et al., 2021), we construct a sketch to illustrate the overall dynamical nature and details of molecular gas flows toward the CMZ. The upper portion of Figure 3 displays an edge-on view of the spatial distribution and motion characteristics of the overall gas flows from the perspective of the observer, while the lower portion shows gas flows dominated by the Galactic bar’s gravitational potential in a face-on view. For the edge-on view, note that the gas inflows extend to in the positive longitude (i.e., the near dust lane DL1; see Su et al., 2024) and in the negative longitude for the far dust lane by assuming that it has a similar length and inclination angle of (between the bar major axis and the Sun-GC) as DL1. The realistic scenario may be much more complex than what we depict here, but this sketch helps us get a comprehensive picture of the entire gas dynamics process toward the CMZ.
@@
Many of the physical processes in the CMZ, including gas dynamics, have been studied in simulations, in which gas is compressed and shocked as a result of the shift from x1 orbit family to x2 orbit family (e.g., Athanassoula, 1992; Englmaier & Gerhard, 1999; Maciejewski et al., 2002; Regan & Teuben, 2004; Sheth et al., 2005; Rodriguez-Fernandez & Combes, 2008; Combes et al., 2014; Sormani et al., 2015). As seen in Figure 3, our observations align well with the model in which the Galactic bar potential (1) channels a substantial amount of molecular gas from the inner Galactic disk at 3.2–3.4 kpc toward the CMZ (i.e., DL1 and the associated kinematic features of CO gas), (2) generates noncircular motions and large-scale shock structures along the near dust lane (e.g., a series of bow shocks revealed by CO gas; Su et al., 2024), (3) governs and regulates the formation and evolution of the CMZ through the overshoot mechanism (i.e., gas flow collision/accumulation near G3/G5/G7 and multiple inflows DL2/DL3 with different trajectories toward the CMZ in Section 2; also see details in the following discussion), and (4) has important impacts on the gas distribution, properties, and subsequent star formation within the inner Galaxy (e.g., the 3 kpc arm, the thickness of the gas disk and the star formation rate therein, etc.). As another good example of gas influenced by the nonaxisymmetric bar potential, we will further discuss the tilted bar lanes traced by CO flows in Section 4.
@@
The Galactic bar systematically transports a massive inflow toward the CMZ. However, this is a necessary but not a sufficient condition for the accumulation of gas in the CMZ. Another key question is how to effectively remove the large amount of angular momentum of the gas located a few kiloparsecs away, so that the gas can settle continuously into the region within the central few hundred parsecs of the Milky Way. Next, we will focus on how the overshoot effect plays a critical role in regulating the redistribution of gas and its angular momentum, as robustly evidenced by several key kinematic characteristics and dynamical effects observed in the new CO flows toward the CMZ.
@@
In Section 2, we undoubtedly confirmed that G3 originates in violent collisions between the inflowing gas of DL2/DL3 and the overshoot gas from the far dust lane (the G3 case for x1 orbits intersects with x2 orbits; Figure 2(b)). In this way, G3 is located near the outskirts of the CMZ (i.e., at 0.5 kpc). Meanwhile, gas flow collisions also occur in regions where the overshoot gas changes its trajectory and passes through the bar channel due to varying torques. When the overshoot gas enters the bar channel, collisions between gas flows are inevitable at the intersections of different eccentric orbits (the G5 and G7 cases for gas flow interactions in bar channel; Figures 2(a) and (c)). Therefore, collisions between gas flows are very common in extremely crowding orbits, especially in the region where the x1 orbits of gas intersect with the x2 orbits in nonaxisymmetric gravitational potential (Figure 3).
@@
The starting points of DL2 and DL3 are located exactly around the regions of G7 and G5, respectively. And the gas mass in the intermediate samples is 1 order of magnitude higher than that in the overshoot gas, confirming significant accumulation of the overshoot gas within the bar channel at these certain locations (e.g., near the G7, G5, and G3 regions). The consistency may indicate that, due to the dynamics of the rotating bar, a significant accumulation of overshoot gas occurs in some special regions (i.e., near G5 at and G7 at by adopting the bar length of 3.2–3.4 kpc; see Su et al., 2024). In addition, considerable CO emission in DL1 is also enhanced significantly near these regions ( and in Figure 2(c); also see Figure 1 in Su et al., 2024). These regions are located near the apocenter of the overshoot flows (Figure 3). The interesting feature needs to be further explored.
@@
Importantly, the measured line-of-sight velocities of the DL2 and DL3 inflows toward the CMZ are 1.5–8 times smaller than those of the DL1 at a similar (see Figure 1; also refer to the geometry proposed in Su et al., 2024). The observations clearly indicate that the tangential velocity of gas rotating clockwise around the GC has significantly decreased, and the orbital angular momentum of overshoot gas flows has been reduced by approximately 0.2–0.7 dex after being captured and reaccelerated by the bar potential. These features are consistent with the fact that there is substantial exchange of mass and angular momentum as overshoot gas changes from approaching flows to receding flows at some locations (e.g., the G7 region at 2.2 kpc, the G5 region at 1.6 kpc, and the G3 region at 0.5 kpc; see Figures 2 and 3). Notably, our observations provide the first clear evidence that the overshoot gas, after being captured by the bar potential during its passage through the bar channel, continues to flow toward the CMZ with significantly reduced angular momentum.
@@
The gas mass in DL1 and that associated with the overshoot effect are and (Table 1), respectively. The result indicates that most of the accretion gas does not enter the CMZ directly but rather overshoots into the bar channel on the other side. Based on our observations, the current inflow efficiency can be easily obtained to be 1/3 on average, which agrees well with the simulation (e.g., Hatchfield et al., 2021). Considering the importance of the overshoot effect (e.g., Maciejewski et al., 2002; Sormani & Barnes, 2019; Hatchfield et al., 2021), we estimate an average inflow rate as . Here, 1/3 is the observed inflow efficiency, while 2 corresponds to both the near and far bars. And 15 Myr comes from the gas kinematics of the DL2 and DL3 inflows by adopting an average inflow velocity of (see DL2 and DL3 in Figure 1). Note that the instantaneous accretion rate can be several times higher than the average value. For example, it can easily cause the accretion rate to increase by nearly five times as several MCs (e.g., with a mass of and a size of pc) with low angular momentum (and thus higher inflow efficiency than 1/3) settle into the CMZ within 0.1 Myr. The gas accumulated in the G3/G5/G7 regions serves as a potential reservoir to regulate the instantaneous gas inflow rate.
@@
Again, not all bar-driven gas flows directly enter the CMZ; instead, they take approximately three orbital periods to settle into the CMZ (i.e., the observed small inflow efficiency of 1/3). Through this process, the gas is subjected to intense gravitational torques from the bar, causing it to be stretched, distorted, shredded, torn apart, and fragmented (Figure 2(c)). Apparently, the overshoot effect greatly lengthens the action time of various dynamical processes before the accretion gas enters the CMZ. Throughout the entire dynamical process, the angular momentum of the gas changes and redistributes due to the combined effects such as torque interactions from the bar, inelastic cloud–cloud collisions, gas viscosity, and various instabilities. The shock dissipation generated during the sufficient dynamics will redistribute angular momentum outward. This enables the distorted gas flows to efficiently transport material from kiloparsec scales to regions within a few hundred parsecs of the GC while also supporting the scenario of the periodic migration of gas toward the CMZ.
@@
Briefly, the confirmation of the overshoot gas, together with frequent collisions caused by intersecting flows, provides detailed observational evidence to explain the dynamics of gas flows toward the CMZ. The overshoot gas with lower angular momentum gradually converges into the bar channel, then flows toward the CMZ again (see DL2 and DL3 in Figure 3). Because of the overshoot effect and frequent collisions, the accretion gas is able to efficiently transfer its excess angular momentum outward through various dynamical processes and then cause the gas to gradually flow toward the CMZ. Here, the gravitational torque of the robust rotating Galactic bar makes a large contribution in the process of redistributing the gas and removing angular momentum from the molecular gas that eventually settles into the CMZ. Due to the interaction of various torques with viscous gas and the frequent collisions between gas flows, at the same time, the strong DISSIPATION process is also essential for the effective outward transfer of gas angular momentum, which in turn remarkably regulates the gas flow dynamics and gas accumulation in the CMZ over the secular evolution of the Galaxy.
@@ 4 Large-scale Tilted Gas Flows and the Twisted Structure in the CMZ
The tilted dust lane has been discussed in many works (e.g., Liszt, 2006, 2008; Rodriguez-Fernandez et al., 2006; Marshall et al., 2008; Sormani & Barnes, 2019; Tress et al., 2020; Akhter et al., 2021). We also confirm that the near dust lane DL1 with a length of –3.4 kpc exhibits distorted features during the large-scale gas accretion driven by the Galactic bar. Additionally, DL1 is distinctly located below the Galactic plane, which is exactly the opposite of the gas flow of the far dust lane above the plane (see Figure 3).
@@
Our previous studies showed that the slightly tilted midplane to the IAU-defined plane (i.e., , see Figure 7 in Su et al., 2016) is caused by the Sun’s height above the physical midplane of the Milky Way (i.e., 15–17 pc; see Su et al., 2016, 2019, 2021). However, the tilted large-scale gas flows toward the CMZ cannot be explained by this (e.g., see discussions in Bland-Hawthorn & Gerhard, 2016) because the maximum deviation of the CO gas structure reaches 170 pc (or at ) from the Galactic plane of . The deviation value is about 2 times larger than the thickness of the thin CO disk (see Table 3 in Su et al., 2021). Considering the smaller thickness of the molecular gas disk in the inner 3–4 kpc region, this deviation from the plane is even more severe. We attribute the large-scale gas flows to a true tilted feature relative to the Galactic midplane. Based on our CO data and model (Su et al., 2024), the tilted angle of the bar lanes relative to the Galactic midplane in is estimated to be , which roughly agrees with the simulations (e.g., in Tress et al., 2020). Indeed, the gas structure revealed by the H i emission exhibits a more extreme tilted characteristic (i.e., or 350 pc for the atomic gas with the most negative velocity; see Burton & Liszt, 1983).
@@
The vertical distribution of gas (e.g., the tilted and warped gas flows) is modulated by the strong gravitational torque generated by the rotating bar and various dissipative forces. In our observation, the gravitational potential of the bar can effectively trap the gas flows into the dust lane (Figure 3). The gas on different orbits is perturbed by the nonaxisymmetric gravitational potential, leading it to undergo noncircular motion around the GC and oscillations relative to the bar. Due to the varying torque exerted by the bar’s potential on the gas, gas flowing along different trajectories and inclinations will move with varying precession rates until it eventually settles into the bar’s potential, aligns with the bar’s major axis, and then rotates around the GC with the bar. These torques also induce misalignments across different parts of the bar lanes. When gas flows on different orbits are subjected to varying torques, they develop distorted or warped structures as observed in CO data. In the process, the angular momentum of the overshoot flows is also transferred to the bar through the dynamic friction. Alternatively, the distortion of the bar lanes caused by resonant trapping of x1 orbits (e.g., Łokas, 2019) may also influence the distribution and motion of the gas. The ongoing interaction between the gas flows and the bar itself accumulates over time, continuously affecting the bar’s structure and motion as it gains mass and angular momentum by exerting torque on the gas.
@@
The above processes are accompanied by complex dynamical interactions that alter the angular momentum and motion states of the gas flows. We confirmed that collisions between gas flows are frequent during the dynamical processes in the G3, G5, and G7 regions. On the one hand, the nonaxisymmetric bar decouples the gas inflows from their initial dynamical state, resulting in rapid assimilation of the accretion gas into the bar potential (e.g., DL2 and DL3). On the other hand, the frequent collisions between gas flows (G3 and G5), combined with the tangential shear of the tidal forces exerted by the nonaxisymmetric bar, result in significant dissipation in the twisted gas flows. Indeed, the typical timescale of these dissipative processes, which are driven by shocks and strong turbulence, may be comparable to the short dynamical timescales associated with gas flow collisions. During the dynamical process of overshoot gas settling into the Galactic bar potential, the dissipation from collisions between misaligned gas flows, together with gas viscosity and the effects of various torques, serves as an efficient mechanism for the outward transfer of angular momentum of gas, thus facilitating their gradual migration toward the CMZ.
@@
The scenario is strongly supported by our observations that the overshoot gas is not fully relaxed into the Galactic bar potential. That is, the overshoot gas from the far dust lane exhibits a more tilted structure (i.e., 1.5 times larger; see from to in Figure 2(c)) than the near dust lane DL1 below the Galactic plane, while a portion of the gas passing through the bar channel even extends above the Galactic plane (i.e., DL2 is above and almost parallel to the Galactic plane). The distorted or warped gas structures in observations may represent the oscillatory characteristics of gas flows perpendicular to the Galactic bar as it moves and relaxes in the bar potential.
@@
Meanwhile, the motion state of the self-consistent bar evolves over time through the exchange of angular momentum between the bar itself and the accreted material. As a result, the bar gains angular momentum from the captured gas and loses it to the halo and the outer stellar disk (e.g., Queiroz et al., 2021; Wylie et al., 2022; Dillamore et al., 2023), leading to a gradual change in the state of the bar (e.g., the morphology, structure, and kinematics; see Li et al., 2016; Chiba et al., 2021; Clarke & Gerhard, 2022; Lucey et al., 2023; Vislosky et al., 2024). We suggest that the nonaxisymmetric bar is essential for regulating the redistribution of accreted gas and its angular momentum transfer. This self-regulating process may stably maintain large-scale gas accretion and gas recirculation over a considerable timescales in the presence of the bar (e.g., 4–8 Gyr, see Bovy et al., 2019; Nepal et al., 2024).
@@
The accumulated gas in the CMZ inherits the information of the gas supplied by the large-scale inflows, which includes the fresh gas from the disk at 3–4 kpc (with higher angular momentum but lower accretion rate) and the recycled gas from the overshoot effect (with lower angular momentum but a higher accretion rate). The moving direction of the large-scale inflows (DL1) is well consistent with the trend of the structure at the far side of the CMZ (see the red solid arrow in the CMZ of Figure 3). That is, gas along the DL1 moves from the negative latitude at toward the positive latitude at , and then back down toward the negative latitude (e.g., see Molinari et al., 2011; Lipman et al., 2024; Veena et al., 2024; Walker et al., 2024; Karoly et al., 2025).
@@
Simultaneously, the inflowing gas along the far dust lane moves from the positive latitude to the negative latitude at , then toward the positive latitude at , and finally back down toward the negative latitude again (i.e., the near side of the CMZ; see the blue solid arrow in the CMZ of Figure 3). Similar to the previous discussion, when the gas flows pass near the x2 orbit, not all gas directly enters the CMZ; instead, some gas experiences the overshoot effect because of its higher angular momentum. Our data have clearly demonstrated that the overshoot gas flew past the near side of the CMZ and then collided violently with the inflowing gas from the DL2 and DL3 (i.e., see the G3 region in Figure 2(b); also see clouds studied by Busch et al., 2022).
@@
We thus propose that the twisted feature of the CMZ (e.g., Molinari et al., 2011; Kruijssen et al., 2015; Ridley et al., 2017) arises from the combined effects of the large-scale tilted gas flows in the dust lanes and the locally distorted gas structure dominated by the strong gravitational potential within a few hundred parsecs of GC (i.e., the nuclear stellar disk; see Sormani et al., 2022). The orbit periods of the twisted structure of the CMZ are about 2–4 Myr (Molinari et al., 2011; Kruijssen et al., 2015), suggesting a rapid precession of the gas under the influence of external nonaxisymmetric gravitational potential and the large-scale tilted gas inflows (e.g., Tress et al., 2020). The timescale of several million years mentioned above is consistent with the dynamical time (2–3 Myr) of gas inflow from the G3 region into the CMZ, indicating that external gas input significantly shapes the geometric configuration and dynamic properties of the CMZ (e.g., local gas accumulation, twisted structure, and rapid precession). Recent observations have revealed that the molecular circumnuclear disks with different kinematics are misaligned with the large-scale gas structure (Combes et al., 2019; Ruffa et al., 2019, 2020; Papachristou et al., 2023), indicating that highly warped gas structures are likely prevalent in the central region of galaxies.